OpenStudy (anonymous):

What is the slope-intercept form equation of the line that passes through (2, 4) and (4, 10)? y = −3x − 2 y = −3x + 2 y = 3x + 2 y = 3x − 2

OpenStudy (anonymous):

well first you would have to find the slopes of the 2 coordinates (2, 4) and (4, 10). Do u know how to find the slope?

OpenStudy (anonymous):

\[slope = m = \frac{ y _{2}- y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

\[m = \frac{ 10 - 4 }{ 4-2 } = \frac{ 6 }{ 2 } = 3\]

OpenStudy (anonymous):

so, the slope is 3. m= 3 in \[y = mx + b\]

OpenStudy (anonymous):

Automatically, the first and second options to the question are out because the slope there is -3 instead of 3.

OpenStudy (anonymous):

Now, you can use any of the two coordinates (2, 4) or (4, 10) to substitute for y and x in order to find b. Let's choose (2, 4).

OpenStudy (anonymous):

\[y = 3x + b\] substitute 2 for x and 4 for y to the above equation because (2, 4) is equal to (x, y) \[4 = 3(2) + b \] \[4 = 6 + b\] solve for b. 4 = 6 + b -6 -6 ---------- b = -2

OpenStudy (anonymous):

Now you found out what m is equal to and what b is equal in order to write the slope-intercept form. So, this is the answer: the last option \[y = 3x -2\]

OpenStudy (anonymous):

thank you!